OpenStudy (anonymous):

Determine whether the infinite geometric series is convergent or divergent. If it is convergent, find its sum. 1 + 1/2 + 1/4 + 1/8 + ...

OpenStudy (anonymous):

Convergent, this is a geometric series with common ratio \(\frac 12 \) The sum is \[ \Huge \frac {1 }{1-\frac 12 } =2\]

OpenStudy (anonymous):

thanks!